OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1 x = 0 x = 2 There is no x value as there is no solution. x can be any value, as there is an infinite number of solutions

OpenStudy (jdoe0001):

so..... if you solve say " 2x + y = 1 " for "y" what would that look like?

OpenStudy (anonymous):

uhm... I been studying this for 3 days now and I don't get it

OpenStudy (jdoe0001):

well.. what if you say subtract 2x from both sides ?

OpenStudy (phi):

they want you to solve for x using substitution 2x + y = 1 4x + 2y = -1 if you use the first equation: 2x +y= 1 add -2x to both sides: 2x - 2x + y = 1 -2x simplify the left side: y= 1-2x that means y can be replace by (1-2x) so do that in the second equation 4x + 2(1-2x) = -1 now "solve". you will find you get a strange result... meaning there is no solution.